# Context-free graphs and their transition groups

### v2

@@ 10. Some examples using the free product of inverse graphs
Let $\chi_{c}:A^{*}\to\mathbb{Z}$ be the homomorphism counting the $c$, i.e., $\chi_{c}(c)=-\chi_{c}(c^{-1})=1$ and deleting all the occurrences of $a$. Note that $\mathfrak{T}$ is locally-quasi-transitive with $\mathfrak{T}\setminus D_{1}(x_{0})=\Lambda_{1}\setminus D_{1}(y_{0})$, where $\Lambda_{1}$ is the transitive graph formed by two $c$-lines with two loops labeled by $a,a^{-1}$ everywhere. By Lemma 7.5, the language $\mathcal{O}_{1}$ is formed by words that label a circuit at every vertex of $\Lambda_{1}$. It is easy to see that this is exactly the set of words $u\in A^{*}$ with $\chi_{c}(u)=0$. Therefore, for any vertex $z_{i}^{(j)}$, $j\in\{0,1\}$ of $\mathfrak{T}$ and $u\in\mathcal{O}_{1}$ we have that either $z_{i}^{(j)}\cdot u=z_{i}^{(j)}$, or $z_{i}^{(j)}\cdot u=z_{i}^{(j+1\mod 2)}$, where the second case occurs if and only the walk $z_{i}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}z_{i}^{(j+1\mod 2)}$ passes from the edges $z_{0}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}z_{0}^{(j+1\mod 2)}$, $x\in\{a,a^{-1}\}$, an odd number of times. In particular we have $z_{i}^{(j)}\cdot u^{2}=z_{i}^{(j)}$ for every vertex of $\mathfrak{T}$, whence every element of $\mathrsfs{B}_{1}$ has torsion of order two. The boundary subgroup $\mathrsfs{B}_{1}$ is infinite. Indeed, take the elements $h_{n}=c^{-n}ac^{n}\in\mathrsfs{B}_{1}$, $n\in\mathbb{Z}$. These are all distinct elements in $\mathcal{G}(\mathfrak{T})$ since for any $j\in\{0,1\}$ we have
@@
We may explicitly characterize the word problem $\mathcal{L}(\mathfrak{T})$. We first need some definitions. For a word $u\in A^{*}$, let us denote by $u[i]$ the letter of $u$ occurring at position $i$, and by $u[:i]$ let us denote the prefix of $u$ of length $u$. We say that an occurrence of the letter $a=u[i]$, or $a^{-1}=u[i]$, is at level $m$ if $\chi_{c}(u[:i])=m$. Since there are just two vertices, namely $z_{0}^{(1)},z_{0}^{(1)}$ in which a walk may enter, it is not difficult to see that for a fixed $m$ the number of $a,a^{-1}$ at level $m$ in $u$ represents the number of times a walk starting from $z_{-m}^{(k)}$, $k\in\{0,1\}$, passes through the “bridge” given by the edges $z_{0}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}z_{0}^{(j+1\mod 2)}$, $x\in\{a,a^{-1}\}$. Therefore, a word $u\in A^{*}$ represents the identity in $\mathcal{G}(\mathfrak{T})$ if and only if $\chi_{c}(u)=0$ and for each $m\in\mathbb{Z}$ the number of $a,a^{-1}$ at level $m$ is even. Note that we clearly have finitely many $m$ with non-zero number of $a,a^{-1}$. We know that this language is CoCF, however, is it also poly-context-free? Let us state this in the following open problem.

### v3

@@ 10. Some examples using the free product of graphs
Let $\chi_{c}:A^{*}\to\mathbb{Z}$ be the homomorphism counting the $c$, i.e., $\chi_{c}(c)=-\chi_{c}(c^{-1})=1$ and deleting all the occurrences of $a$. Note that $\mathfrak{T}$ is locally-quasi-transitive with $\mathfrak{T}\setminus D_{1}(x_{0})=\Lambda_{1}\setminus D_{1}(y_{0})$, where $\Lambda_{1}$ is the transitive graph formed by two $c$-lines with two loops labeled by $a,a^{-1}$ everywhere. By Lemma 8.5, the language $\mathcal{O}_{1}$ is formed by words that label a circuit at every vertex of $\Lambda_{1}$. It is easy to see that this is exactly the set of words $u\in A^{*}$ with $\chi_{c}(u)=0$. Therefore, for any vertex $z_{i}^{(j)}$, $j\in\{0,1\}$ of $\mathfrak{T}$ and $u\in\mathcal{O}_{1}$ we have that either $z_{i}^{(j)}\cdot u=z_{i}^{(j)}$, or $z_{i}^{(j)}\cdot u=z_{i}^{(j+1\mod 2)}$, where the second case occurs if and only the walk $z_{i}^{(j)}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}z_{i}^{(j+1\mod 2)}$ passes from the edges $z_{0}^{(j)}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$x$}}}{{\relbar\joinrel\longrightarrow}}}z_{0}^{(j+1\mod 2)}$, $x\in\{a,a^{-1}\}$, an odd number of times. In particular we have $z_{i}^{(j)}\cdot u^{2}=z_{i}^{(j)}$ for every vertex of $\mathfrak{T}$, whence every element of $\mathrsfs{B}_{1}$ has torsion of order two. The boundary subgroup $\mathrsfs{B}_{1}$ is infinite. Indeed, take the elements $h_{n}=c^{-n}ac^{n}\in\mathrsfs{B}_{1}$, $n\in\mathbb{Z}$. These are all distinct elements in $\mathcal{G}(\mathfrak{T})$ since for any $j\in\{0,1\}$ we have
@@
We may explicitly characterize the word problem $\mathcal{L}(\mathfrak{T})$. We first need some definitions. For a word $u\in A^{*}$, let us denote by $u[i]$ the letter of $u$ occurring at position $i$, and by $u[:i]$ let us denote the prefix of $u$ of length $u$. We say that an occurrence of the letter $a=u[i]$, or $a^{-1}=u[i]$, is at level $m$ if $\chi_{c}(u[:i])=m$. Since there are just two vertices, namely $z_{0}^{(1)},z_{0}^{(1)}$ in which a walk may enter, it is not difficult to see that for a fixed $m$ the number of $a,a^{-1}$ at level $m$ in $u$ represents the number of times a walk starting from $z_{-m}^{(k)}$, $k\in\{0,1\}$, passes through the “bridge” given by the edges $z_{0}^{(j)}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$x$}}}{{\relbar\joinrel\longrightarrow}}}z_{0}^{(j+1\mod 2)}$, $x\in\{a,a^{-1}\}$. Therefore, a word $u\in A^{*}$ represents the identity in $\mathcal{G}(\mathfrak{T})$ if and only if $\chi_{c}(u)=0$ and for each $m\in\mathbb{Z}$ the number of $a,a^{-1}$ at level $m$ is even. Note that we clearly have finitely many $m$ with non-zero number of $a,a^{-1}$. We know that this language is CoCF, however, is it also poly-context-free? Let us state this in the following open problem.
@@
Acknowledgments
The authors are members of the Gruppo Nazionale per le Strutture Algebriche, Geometriche e le loro Applicazioni (GNSAGA) of the Istituto Nazionale di Alta Matematica (INdAM).
The second author is also a member of the PRIN 2022 “Group theory and its applications” research group of the Ministero dell’università e della ricerca and gratefully acknowledges its support.
The third author acknowledges support from the Grant QUALIFICA by Junta de Andalucía grant number QUAL21 005 USE and from the research grant PID2022-138719NA-I00 (Proyectos de Generación de Conocimiento 2022) financed by the Spanish Ministry of Science and Innovation.
We are deeply grateful to Corentin Bodart for helpful discussions on the first draft of this manuscript, especially for Proposition 4.4, Proposition 10.5 and for Section 5.